OpenStudy (anonymous):

Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(one-third) is undefined. Be sure to show your work.

OpenStudy (anonymous):

Will fan!

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Is there a function that comes with this problem or do we need to make up a function?

OpenStudy (anonymous):

well the first part of the question is : Simplify f(x) = the quantity 3 x-squared minus 4x plus 1 all over 3x minus 1. but idk if that goes wit it

OpenStudy (anonymous):



OpenStudy (anonymous):

with*

OpenStudy (anonymous):

@mathstudent55 @Paynesdad

OpenStudy (mathstudent55):

It does go with it.

OpenStudy (anonymous):

Oh..

OpenStudy (mathstudent55):

\( f(x) = \dfrac{3x^2 -4x+1}{3x - 1} \) L:et's find f(1): \( f(1) = \dfrac{3(1^2) -4(1)+1}{3(1) - 1} = \dfrac{3 -4+1}{3 - 1} = \dfrac{0}{2} = 0\) Ok?

OpenStudy (anonymous):

okay!

OpenStudy (mathstudent55):

You see that as the problem read, f(1) = 0. That make sense.

OpenStudy (anonymous):

Yes so do I just fill in x with what ever f is ? Also how do you know if something is underfined?

OpenStudy (mathstudent55):

Now let's calculate f(0) \(f(0) = \dfrac{3(0^2) -4(0)+1}{3(0) - 1} = \dfrac{0 -0+1}{0 - 1} = \dfrac{1}{-1} = -1\) Once again, we confirm that the given info is correct. f(0) = -1

OpenStudy (mathstudent55):

Now we can calculate f(-1): \(f(-1) = \dfrac{3(-1)^2 -4(-1)+1}{3(-1) - 1} = \dfrac{3 +4+1}{-3 - 1} = \dfrac{8}{-4} = -2\) Once again, we have shown the given info to be true. f(-1) = -2.

OpenStudy (mathstudent55):

Now comes the last step, to show that \(f \left( \dfrac{1}{3} \right) \) is undefined. Usually in functions, you get undefined results when they involve division by zero.

OpenStudy (mathstudent55):

\(f(\frac{1}{3}) = \dfrac{3(\frac{1}{3})^2 -4(\frac{1}{3})+1}{3(\frac{1}{3}) - 1} \) Can you simplify the denominator? What do you get for the denominator?

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

*

OpenStudy (mathstudent55):

No. \( 3(\dfrac{1}{3}) - 1 = ? \) What is 3 times one-third?

OpenStudy (anonymous):

oh thought it was 1/2 not 1/3 and its zero so i getttt ittt!!! THanks sooo much!

OpenStudy (mathstudent55):

Correct. Since the denominator is zero when x = 1/3, the function is undefined at x = 1/3.